INAF queue observing, 2015 Nov 04-Nov 05

Observer: S. Bisogni, R. Carini
Telescope Operator: Steve Allanson
Support Astronomer: Barry Rothberg
Instrument: LUCI

Summary

00:41 UT the humidity is 96% and the dome is completely covered by snow. Tonight it snowed very hard and we are waiting for Santa to come. The staff informed us that for tonight there is not chance at all to open.

Night duration (12\xB0 el twilight) 19:00-5:10 ( 02:00-12:10 UT) = 10 h 10m

Observing time: 0 hrs 0 m

Weather time loss: 10 h 10 m (100%)

technical time loss 0 m

Data Summary

Project_RUN Objects Filter ExpTime Notes
Calibrazioni Ragazzoni 942707_PY_315      
Calibrazioni Ragazzoni 951655.PY_290      

Details

The Astronomer Support makes test with MODS2 and LUCI2
06:35 we make test for the calibrations for Iovino and Ragazzoni programs to find the right exposure time for flat field and arc. luci.2015.1105.0037-0051 10:35 star the calibration for Ragazzoni 951655.PY_290 luci.2015.1105.0052-0071==flat, ==luci2.0074-0075,luci.0078-0079,luci.0082-0083 arc the others are images of persistence.
  • 942707_PY_315 luci.2015.1105.0084-00103==flat, are not good ==luci2.0106-0107,luci2.0110-0111,luci2.0114-0115 arc, the other are of persistence.
Article text.

-- %USERSIG{SusannaBisogni - 2015-11-05}%

Comments

 
Topic revision: r5 - 07 Nov 2015, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback